Linkin Corporation is considering purchasing a new delivery truck. The truck has many advantages over the company’s current truck (not the least of which is that it runs). The new truck would cost $56,760. Because of the increased capacity, reduced maintenance costs, and increased fuel economy, the new truck is expected to generate cost savings of $8,600. At the end of 8 years, the company will sell the truck for an estimated $28,600. Traditionally the company has used a rule of thumb that a proposal should not be accepted unless it has a payback period that is less than 50% of the asset’s estimated useful life. Larry Newton, a new manager, has suggested that the company should not rely solely on the payback approach, but should also employ the net present value method when evaluating new projects. The company’s cost of capital is 8%.

Click here to view the factor table.

(a)

Compute the cash payback period and net present value of the proposed investment. (If the net present value is negative, use either a negative sign preceding the number eg -45 or parentheses eg (45). Round answer for present value to 0 decimal places, e.g. 125. Round answer for Payback period to 1 decimal place, e.g. 10.5. For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

Cash payback period Enter a number of years rounded to 1 decimal place  years
Net present value $ enter a dollar amount rounded to 0 decimal places


(b)

Does the project meet the company’s cash payback criteria?

select an option                                                          NoYes


Does it meet the net present value criteria for acceptance?

select an option                                                          YesNo
Answer #1

Answer:

a) Cash payback period = Initial investment/Annual cash flow

= 56760/8600

Cash payback period = 6.6 years

Net present value = Present value of cash inflow-Present value of cash outflow

= (8600*5.74664+28600*0.54027)-56760
= 49421.104+15451.722-56760

Net present value = 8113

b) No project does not meet the company's cash payback criteria

Yes, it does meet the net present value criterial for acceptance